Please answer correctly thank you.
Andre KArana . Check My Weork Exercise 84) had a driers leense. Twenty-ve years terr thet percentage had dropped to TSS (University of Michigan Transportation Research Institute website, April 7,2012), Suppose these results are based on are based on a random ample of L200 ǐ9-yearalds n i 98, and again in 2006. estimate of the number of 19-yean-old dnvers in 19837 Round your intermedate answers to four deaimal places (four decimal places) Interval estimace (to four decimal places) b.At95% condence, what is the mergn oferror nd the Margin of eror nterval c. Is the mangin of error the same in parts (a) and (b)? estimate of the number of 19-year-old drivers in 20087 Round you Why, or why o The input in the below will not be graded, but may be reviewed and considered by your instrcto Check My Work
Answer #1

1. Margin of error = 1.960*\sqrt{\frac{0.87*0.13}{1200}}

= 0.0190

Interval estimate

0.87-0.0190, 0.87+0.0190

0.8510 to 0.8890

2) Margin of error = 1.960*\sqrt{\frac{0.75*0.25}{1200}}

= 0.0245

Interval estimate

0.75-0.0190, 0.75+0.0190

0.8455 to 0.8945

c) Margin of error are not same because of different proportions